# The Discovery of Little Red Dots in the Local Universe: Signatures of Cool Gas Envelopes

### v1

@@ 3.1 Photometry
The three targets were observed by multiple wide-field sky surveys. We adopt photometric data from the following surveys: FUV and NUV from GALEX DR6 (L. Bianchi & B. Shiao, 2020), ugriz Petrosian photometry from SDSS DR17 (K. N. Abazajian et al., 2009; Abdurro’uf et al., 2022), PSF photometry from Pan-STARRS DR2 (H. A. Flewelling et al., 2020), mean photometry from GAIA DR3 (Gaia Collaboration et al., 2023), photometry from Legacy Surveys DR10 (A. Dey et al., 2019), Petrosian photometry from UKIDSS DR11PLUS (A. Lawrence et al., 2007), and W1, W2, W3, W4 photometry from WISE (E. L. Wright et al., 2010). The GALEX photometry is taken from the GUVmatch catalog. The WISE photometry has been incorporated into the Legacy Survey catalog based on the unWISE images (E. F. Schlafly et al., 2019). All other data are retrieved from Astro Data Lab444https://datalab.noirlab.edu/ (M. J. Fitzpatrick et al., 2014; R. Nikutta et al., 2020).
3.2 Spectroscopic Follow-up Observations
@@
We obtained high signal-to-noise-ratio (S/N) LBT/MODS and Magellan/FIRE spectroscopy to characterize the UV-to-NIR spectra of the three local LRDs. We also obtained MMT/Binospec spectra of J1025+1402 and J1022+0841 to measure their high-resolution H profiles.
@@
The LBT/MODS (R. W. Pogge et al., 2010) is a pair of dual-channel spectrographs on the 28.4m Large Binocular Telescope, with the blue channel covering 3300–6000 Å and the red channel 5000–10000 Å. We observed the three targets on February 24, 2025, using 1.0-arcsec slits in grating mode under an average seeing of 0.8 arcsec. This setup achieved a spectral resolution of –1500 in the blue channel and –1700 in the red channel. Observations of each target included 3600 s exposures in both the blue and red channels, with the two LBT telescopes observing simultaneously. As a result, each target effectively has 6600 s exposures in both the blue and red channels.
@@
The Magellan/FIRE (R. A. Simcoe et al., 2013) is a near-IR echelle spectrograph covering 0.8–2.5 on the 6.5m Magellan Baade Telescope. We observed J1025+1402 using a 0.75-arcsec slit on December 21, 2024, reaching –6000. The observation included 5900 s exposures. We observed J1047+0739 and J1022+0841 using 1-arcsec slits on March 11, 2025, with 7900 s exposures for J1047+0739 and 6900 s exposures for J1022+0841, reaching –4000. The seeing for these observations was –0.7 arcsec.
@@
The MMT/Binospec observation of J1025+1402 was conducted in the IFU mode (D. Fabricant et al., 2019, 2025) on December 7, 2024, using the 600 line/mm grating. The observation consisted of s exposures and achieved a spectral resolution of . The seeing during the observation was approximately 1.6 arcsec. For J1022+0841, the observation was conducted in long-slit mode on January 23, 2024, using a 1-arcsec slit and the 600 line/mm grating, under a seeing of approximately 2.3 arcsec. The observation included s exposures, with a spectral resolution of .
@@
The LBT/MODS and Magellan/FIRE data were reduced with PypeIt (J. Prochaska et al., 2020). We performed bias subtraction, flat-fielding, wavelength calibration, sky subtraction, spectral extraction, and telluric correction. To correct for slit losses and calibrate the absolute flux of the LBT/MODS spectra, we first matched them to the SDSS spectra using a wavelength-dependent linear function. The spectra in SDSS DR17 have been processed with an updated flux calibration algorithm, yielding significant improvements at wavelengths below 6000 Å555https://www.sdss4.org/dr17/algorithms/spectrophotometry/. We confirmed that the SDSS spectra closely follow the shape of the corresponding photometry. Then, we applied a constant scaling factor to align the semi-calibrated spectra with the SDSS photometry. The Magellan/FIRE spectra were scaled to match the UKIDSS -band photometry using a constant factor. The MMT/Binospec data were reduced using the Binospec pipeline (J. Kansky et al., 2019). The IFU data were extracted with a rectangular aperture matching the seeing disk. Throughout this work, most measurements are based on MODS and FIRE spectra, while Binospec spectra are used exclusively to analyze the H absorber profiles of J1025+1402 and J1022+0841.
@@
Figure 1 shows the UV-to-NIR spectra, multi-wavelength photometry, and images of the three targets. The LBT/MODS spectra confirm the presence of the V-shape of the UV-to-optical continua. The declining near-IR spectra in space are consistent with the rest-frame IR photometry of bright high-redshift LRDs detected by MIRI (e.g., B. Wang et al., 2025a; A. de Graaff et al., 2025; A. J. Taylor et al., 2025a; D. J. Setton et al., 2025).
@@ 4 Evidence as local LRDs
Basic property
Name | RA | DEC | (mag) | 12 + | ||||
( erg s-1) | ( erg s-1) | ( erg s-1) | ||||||
J102530.29+140207.3 | 156.37622 | 14.03586 | 0.1007 | 19.003 | 7.47 0.30 | 4.95 | 7.550.03 | |
J104755.92+073951.2 | 161.98302 | 7.66423 | 0.1682 | 18.846 | 18.5 0.53 | 8.000.02 | ||
J102208.52+084156.1 | 155.53552 | 8.69892 | 0.2227 | 20.011 | 17.4 0.31 | – | 7.430.03 |
H emission-line and BH properties
Name | FWHM | FWHMHα,narrow | |||
( erg s-1) | ( erg s-1) | (km s-1) | (km s-1) | ||
J1025+1402 | 8.68 0.13 | 1.33 0.02 | 983 10 | unresolved | |
J1047+0739 | 27.51 0.58 | 18.86 0.15 | 949 25 | 47 3 | 6.50 0.09 |
J1022+0841 | 48.10 0.49 | 3.43 1.10 | 1022 18 | 6.70 0.09 | |
@@
In this section, we present the observed characteristics of three local objects that closely resemble high-redshift LRDs. The section is organized as follows: morphology and continuum luminosities are discussed in §4.1; broad emission lines and BH masses in §4.2; Balmer and He I absorption features in §4.3; variability properties in §4.4; X-ray and radio properties in §4.5; and a summary of the observational evidence as LRDs is provided in §4.6. We present their zoomed-in spectra in Figures 2, 3, and 4, and summarize the derived properties in Table 1.
4.1 Morphology and Continuum Luminosity
@@
All three objects exhibit very compact morphologies across the optical bands. They are not spatially resolved in LBT/MODS direct imaging with a seeing of 0.9″. This implies that the diameter of J1025+1402 is comparable to or smaller than 1.7 pkpc, while those of J1047+0739 and J1022+0841 are constrained to be smaller than 2.5 pkpc and 2.3 pkpc, respectively.
@@
J1047+0739 was observed with the HST/ACS FR782N ramp filter narrow-band imaging that covers both the H line and the adjacent continuum (GO-15617; PI: F. Bauer), achieving a spatial resolution of 206 pc. The continuum map reveals a bright central point source contributing about 85% of the light, alongside a compact yet clumpy host galaxy with a half-light radius of 530 pc. The H line map is dominated by the central point source with minimal host contribution. The spatially resolved map shows that, in J1047+0739, AGN emission dominates both the continuum and the H line, in agreement with our definition of LRDs, although with a non-negligible contribution of the stellar light from the host galaxy to the continuum. The detailed morphological analysis is presented in Appendix A. Future high-resolution imaging with HST, JWST, or ground-based AO will further constrain the properties of the physical sizes of the local LRDs and their host galaxy properties.
@@
All three objects exhibit continuum shapes characterized by pronounced V-shapes and declining IR slopes. The inflections occur near the Balmer limit. The continuum luminosity at rest-frame 5100Å () is around erg s-1. J1047+0739 exhibits a potential Balmer jump, presumably due to the presence of a host galaxy. For all three objects, the WISE photometry at rest-frame 3–20 µm reveals the presence of weak dust emission, at a wavelength range beyond what is possible with JWST observations of high-redshift LRDs. We integrate the continua from rest-frame 2000 Å to 20 µm to estimate the total continuum luminosity over this wavelength range, yielding – erg s-1 for all three objects.
4.2 Broad Emission Lines and BH masses
@@
The three objects show clear broad profiles in H, H, Pa to Pa, and several He I lines (e.g., He I 7065, 10830). J1025+1402 and J1022+0841 also exhibit broad infrared O I emission lines, consistent with the line broadening originating from gas motions in the BLRs around BHs. The methodology of the line measurements is detailed in Appendix B. Specifically for the H lines, we model the narrow component using a single narrow Gaussian. The broad components are fitted with two to three broad Gaussians. The [N II] 6548, 6583 doublet is modeled with two narrow Gaussians sharing the same FWHM as the narrow H component and fixed line ratios of 3. For J1025+1402 and J1022+0841, additional Gaussian absorbers are applied atop the broad emission lines. The measurements from the MODS spectra are summarized in Table 1. We also report the narrow H FWHMs of J1025+1402 and J1022+0841 from the Binospec spectra, taking advantage of its higher spectral resolution.
@@
All three objects exhibit broad H emission with luminosities around erg s-1, and their individual broad components exhibit FWHMs ranging from km s-1 to 2000 km s-1 , with the composite FWHMs around 1000 km s-1. We estimate their BH masses using the H luminosity (), the composite broad-line H FWHM, and the empirical relation from J. E. Greene & L. C. Ho (2005). We find BH masses around –. This scaling relation for BH mass estimates is widely used in high-redshift LRD studies (e.g., J. Matthee et al., 2024; X. Lin et al., 2024, 2025; A. J. Taylor et al., 2025b). However, this estimator might not be fully applicable to LRDs, if their physical nature is different from classical AGNs, particularly if their line profiles and continua arise from distinct physical mechanisms (see §7). The measured values in these objects are 3–6 times higher than the predicted by the – relation in J. E. Greene & L. C. Ho (2005), where refers to the total (broad + narrow) H luminosity from AGNs. It remains unclear whether or reliably traces the BLR size in these systems.
@@ 4.3 Balmer and He I Absorption
Figure 5 presents the line profiles of H, H, He I, and Pa for the three local LRDs. All three objects exhibit absorption lines in their Balmer series or He I 10830 lines. For H, we model the absorber as Gaussian components superimposed on the intrinsic broad H emission. We then simultaneously fit the absorbers, the narrow H, broad H, and the [N II] 6548, 6583 doublet. For H and He I 10830, absorbed regions are masked, and the intrinsic emission line profile is modeled using a narrow Gaussian combined with multiple broad Gaussians. The detailed fitting methodology and complete absorption measurement information are described in Appendix B.
J1025+1402 exhibits a redshifted H absorption feature superimposed on the broad H emission and continuum, with km s-1 and Å. A similar redshifted H absorber is also seen in the triply imaged LRD in the Abell 2744 field (X. Ji et al., 2025). J1025+1402 and J1047+0739 both have H absorption, superimposed on the broad H emission. The H absorber of J1025+1402 has an EW of Å, and is blueshifted by km s-1 relative to the center of the broad H component. The H absorber in J1022+0841 can be described with two Gaussian components, one blueshifted by km s-1 and one redshifted by km s-1. Their EWs are Å and Å, respectively, yielding a total EW of Å. An alternative solution is one Gaussian absorber with km s-1 and EW of Å (see Appendix B). J1047+0739 does not show clear Balmer absorption.
@@
All three targets show prominent blueshifted He I 10830 absorption. The He I 10830 absorber of J1025+1402 and J1047+0739 has multiple components with distinct kinematics (Figure 6). The line profile of the J1025+1402 He I 10830 absorber, as shown in Figure 5, can be described using three Gaussian components, blueshifted with , and km s-1, respectively. This profile may alternatively result from a mixture of saturated He I absorbers. The line profile of the J1047+0739 He I 10830 absorber can be described using five Gaussian components with , and km s-1, respectively. For J1022+0841, the He I 10830 absorber is well described by a single Gaussian component with km s-1. They highlight the prevalence of He I absorption in local LRDs. As shown in Figure 6, for J1025+1402 and J1022+0841, the velocities of the He I and H absorbers are not well aligned, indicating complex outflow kinematics.
For reference, we also show the Pa line profile in Figure 5. No absorption features are associated with Pa. This suggests that the Balmer absorbers originate from hydrogen atoms with a population elevated to the energy level, but not to the level.

### v2

@@ 3.1 Photometry
The three targets have been observed in multiple wide-field sky surveys. We compile photometric measurements from the following datasets, adopting PSF photometry where available since the targets are unresolved in the images: FUV and NUV from GALEX DR6 (Bianchi2020); ugriz PSF photometry from SDSS DR17 (Abazajian2009; Abdurrouf2022); PSF photometry from Pan-STARRS DR2 (Flewelling2020); mean photometry from GAIA DR3 (Gaia2023); model photometry from Legacy Surveys DR10 (Dey2019) where PSF models are used for the three objects; PSF photometry from UKIDSS DR11PLUS (Lawrence2007); and W1, W2, W3, W4 photometry from WISE (Wright2010). The GALEX photometry is taken from the GUVmatch catalog. The WISE photometry has been incorporated into the Legacy Survey catalog based on the unWISE images (Schlafly2019). All other data are retrieved from Astro Data Lab333https://datalab.noirlab.edu/ (Fitzpatrick_SPIE; NIKUTTA2020100411).
3.2 Spectroscopic Follow-up Observations and Data Reduction
@@
We obtained high S/N LBT/MODS and Magellan/FIRE spectroscopy to characterize the UV-to-near-IR spectra of the three local LRDs. We also obtained MMT/Binospec spectra of J1025+1402 and J1022+0841 to measure their high-resolution H profiles.
@@
The LBT/MODS (Pogge2010) is a pair of dual-channel spectrographs on the 28.4 m Large Binocular Telescope, with the blue channel covering 3300–6000 Å and the red channel 5000–10,000 Å. We observed the three targets on February 24, 2025, using 1.0-arcsec slits in grating mode under an average seeing of 0.8 arcsec. This setup achieved a spectral resolution of –1500 in the blue channel and –1700 in the red channel. Observations of each target included 3600 s exposures in both the blue and red channels, with the two LBT telescopes observing simultaneously. As a result, each target effectively has 6600 s exposures in both the blue and red channels.
@@
The Magellan/FIRE (Simcoe2013) is a near-IR echelle spectrograph covering 0.8–2.5 on the 6.5 m Magellan Baade Telescope. We observed J1025+1402 using a 0.75-arcsec slit on December 21, 2024, reaching –6000. The observation included 5900 s exposures. We observed J1047+0739 and J1022+0841 using 1-arcsec slits on March 11, 2025, with 7900 s exposures for J1047+0739 and 6900 s exposures for J1022+0841, reaching –4000. The seeing for these observations was –0.7 arcsec.
@@
The MMT/Binospec observation of J1025+1402 was conducted in the IFU mode (Fabricant2019; Fabricant2025) on December 7, 2024, using the 600 line/mm grating. The observation consisted of s exposures and achieved a spectral resolution of . The seeing during the observation was approximately 1.6 arcsec. For J1022+0841, the observation was conducted in long-slit mode on January 23, 2024, using a 1-arcsec slit and the 600 line/mm grating, under a seeing of approximately 2.3 arcsec. The observation included s exposures, with a spectral resolution of .
@@
The LBT/MODS and Magellan/FIRE data were reduced with PypeIt (Prochaska2020). We performed bias subtraction, flat-fielding, wavelength calibration, sky subtraction, spectral extraction, and telluric correction. To correct for slit losses and calibrate the absolute flux of the LBT/MODS spectra, we first matched them to the SDSS spectra using a wavelength-dependent linear function. The spectra in SDSS DR17 have been processed with an updated flux calibration algorithm, yielding significant improvements at wavelengths below 6000 Å444https://www.sdss4.org/dr17/algorithms/spectrophotometry/. We confirmed that the SDSS spectra closely follow the shape of the corresponding photometry. Then, we applied a constant scaling factor to align the semi-calibrated spectra with the SDSS photometry. The Magellan/FIRE spectra were scaled to match the UKIDSS -band photometry using a constant factor. The MMT/Binospec data were reduced using the Binospec pipeline (Kansky2019). The IFU data were extracted with a rectangular aperture matching the seeing disk. The Binospec spectra of J1025+1402 and J1022+0841 were flux-calibrated by scaling to match the calibrated LBT/MODS spectra after degrading them to the MODS resolution.
@@
Figure 1 shows the UV-to-near-IR spectra, multi-wavelength photometry, and images of the three targets. The LBT/MODS spectra confirm the presence of the V-shape of the UV-to-optical continua. The declining near-IR spectra in space are consistent with the rest-frame IR photometry of bright high-redshift LRDs detected by MIRI (e.g., Wang2025; deGraaff2025; Taylor2025; Setton2025).
@@ 4 Evidence as local LRDs
| Name | J102530.29+140207.3 | J104755.92+073951.2 | J102208.52+084156.1 |
| RA | 156.37622 | 161.98302 | 155.53552 |
| DEC | 14.03586 | 7.66423 | 8.69892 |
| 0.1007 | 0.1682 | 0.2227 | |
| (mag) | 19.003 | 18.846 | 20.011 |
| ( erg s-1) | 7.57 0.31 | 18.67 0.53 | 17.78 0.31 |
| ( erg s-1) | 4.95 | 29.930.01 | 17.72 |
| ( erg s-1) | – | ||
| 12 + | 7.58 0.03 | 8.01 0.02 | 7.43 0.03 |
| (kpc) | 0.568 | ||
| Inflection point (Å) |
| Name | FWHM | FWHMHα,narrow | b | b | |||
| ( erg s-1) | (km s-1) | ( erg s-1) | (km s-1) | () | (erg s-1) | ||
| J1025+1402 | |||||||
| J1047+0739 | |||||||
| J1022+0841 |
@@
In this section, we present the observed characteristics of three local objects that closely resemble high-redshift LRDs. The section is organized as follows: morphology and continuum luminosities are discussed in §4.1; broad H emission line and BH properties in §4.2; Balmer and He I absorption features in §4.3; variability properties in §4.4; X-ray and radio properties in §4.5; and a summary of the observational evidence as LRDs is provided in §4.6. We present their zoomed-in spectra in Figures 2, 3, and 4, and summarize their properties in Table 1 and 2.
4.1 Morphology and Continuum
@@
All three objects exhibit very compact morphologies across the optical bands. They are not spatially resolved in LBT/MODS direct imaging with a seeing of 0.9″. This implies that the diameter of J1025+1402 is comparable to or smaller than 1.7 pkpc, while those of J1047+0739 and J1022+0841 are constrained to be smaller than 2.6 pkpc and 3.2 pkpc, respectively.
@@
J1047+0739 was observed with the HST/ACS FR782N ramp filter narrow-band imaging that covers both the H line and the adjacent continuum (GO-15617; PI: F. Bauer), achieving a spatial resolution of 206 pc. The continuum map reveals a bright central point source contributing about 85% of the light, alongside a compact yet clumpy host galaxy with a half-light radius of 568 pc (see Appendix A). The H line map is dominated by the central point source with minimal host contribution. The spatially resolved map shows that, in J1047+0739, AGN emission dominates both the continuum and the H line, in agreement with our definition of LRDs, although with a non-negligible contribution of the stellar light from the host galaxy to the continuum. The detailed morphological analysis is presented in Appendix A. Future high-resolution imaging with HST, JWST, or ground-based AO will further constrain the properties of the physical sizes of the local LRDs and their host galaxy properties.
@@
All three objects exhibit continuum shapes characterized by pronounced V-shapes and declining IR slopes. We determine the inflection points by fitting a two-component power law to a 300 Å segment of the continuum centered on the minimum within the rest-frame 3000–4500 Å range. The inflections occur near the Balmer limit. We measure the UV continuum slopes at rest-frame 2500 Å, , by fitting a power law to the GALEX/NUV photometry and MODS spectra below 3000 Å. All the three objects have . The optical continuum slopes () are measured from the MODS spectra around rest-frame 4500 Å, after masking emission lines and potential absorption features. J1025+1402 and J1022+0841 exhibit steep , while J1047+0739 shows a flatter, yet still red, optical continuum slope (). J1047+0739 also exhibits a potential Balmer jump, presumably due to the presence of a host galaxy. The continuum luminosity of the three objects at rest-frame 5100 Å () is around erg s-1. For all three objects, the WISE photometry at rest-frame 3–20 µm reveals the presence of weak dust emission, at a wavelength range beyond what is possible with JWST observations of high-redshift LRDs. We integrate the continua from rest-frame 2000 Å to 20 µm to estimate the total continuum luminosity over this wavelength range, yielding – erg s-1 for all three objects.
4.2 Broad H Emission Lines and BH properties
@@
The three objects show clear broad profiles in H, H, Pa to Pa, and several He I lines (e.g., He I 7067, 10833). J1025+1402 and J1022+0841 also exhibit broad infrared O I 1.29 µm emission lines, consistent with the line broadening originating from gas motions in the BLRs around BHs. The methodology for the line measurements is described in Appendix B. For the H lines, the narrow component is modeled with a single Gaussian, while the broad component is fitted with two to three broad Gaussians. The [N II] 6550, 6585 doublet is modeled with two narrow Gaussians sharing the same FWHM as the narrow H component and a fixed line ratio of 3. For J1025+1402 and J1022+0841, we adopt the MMT/Binospec H profile measurements as fiducial and confirm their consistency with the MODS spectra. The resulting H measurements are summarized in Table 2. The effective FWHM of the broad H profiles (FWHM), used to estimate BH masses, is derived from the summed profile of all broad components.
@@
All three objects exhibit broad H emission with luminosities around erg s-1, and their individual broad components exhibit FWHMs ranging from km s-1 to 2000 km s-1 , with the effective FWHMs around 1000 km s-1. We estimate their BH masses using the broad H luminosity (), the effective broad H FWHM (FWHM), and the empirical relation from Reines2015. We find BH masses around –. This scaling relation for BH mass estimates is widely used in high-redshift LRD studies (e.g., Matthee2024; Lin2024; Lin2025; Taylor2025a). If estimated using the Greene2005 relation, the values would be 0.15–0.33 dex lower. However, these estimators might not be fully applicable to LRDs, if their physical nature is different from classical AGNs, particularly if their line profiles and continua arise from distinct physical mechanisms (see §7). The measured values in these objects are 3–7 times higher than the predicted by the – relation in Greene2005, where refers to the total (broad + narrow) H luminosity from AGNs. If using and FWHMHα,broad to estimate following the relations in Greene2005, the resulting values are 0.5–0.7 dex lower than those listed in Table 2. It remains unclear whether or reliably traces the BLR size in these systems. On the other hand, if the broadening of H lines is partly due to electron scattering in dense ionized gas rather than purely from BLR gas motion, the derived could be overestimated by 1–2 dex (Naidu2025; Rusakov2025). However, whether this scenario applies to the general LRD population remains uncertain (Juodzbalis2025; Brazzini2025).
We note that no dust attenuation correction has been applied in the estimates above. The intrinsic Balmer decrements in BLRs remain uncertain because of the complex radiative transfer effects in high-density gas clouds. The broad H/H line ratios in the three objects are 8-34 (see §5.1 for a more detailed analysis). Assuming Case B recombination, although it may not hold in BLRs, such large ratios correspond to under an SMC attenuation curve (Pei1992). If this attenuation correction is applied, the intrinsic H luminosity could be a factor of higher than the directly measured values, implying that the BH masses would increase by dex. However, if energy balance holds, such a high level of dust attenuation is in tension with the weak IR emission observed by WISE.
For reference and comparison with literature LRDs, we also list bolometric luminosities derived solely from H following Stern2012, along with the corresponding Eddington ratios, in Table 2 ( and ). These values are generally consistent with high-redshift LRDs, which span a wide range of from to , and from to (e.g., Maiolino2024; Greene2024; Lin2024; Juodzbalis2025). We emphasize, however, that these -derived estimates may not accurately reflect the intrinsic bolometric luminosities and Eddington ratios due to the uncertain physical processes in such systems and the systematics discussed above (Greene2025).
@@ 4.3 Balmer and He I Absorption
Figure 5 presents the line profiles of H, H, He I, and Pa for the three local LRDs. All three objects exhibit absorption lines in their Balmer series and/or He I 10833 lines. For H of J1025+1402 and J1022+0841, we model the absorber as Gaussian components superimposed on the intrinsic broad H emission. We then simultaneously fit the absorbers, the narrow H, broad H, and the [N II] 6550, 6585 doublet. For H of J1025+1402 and He I 10833 of all three objects, absorbed regions are masked, and the intrinsic emission line profile is modeled using a narrow Gaussian combined with multiple broad Gaussians. The detailed fitting methodology is described in Appendix B. The measured absorption properties are presented in Table 3.
J1025+1402 exhibits a redshifted H absorption feature superimposed on the broad H emission and continuum, with km s-1 and Å. A similar redshifted H absorber is also seen in the triply imaged LRD in the Abell 2744 field (Ji2025). J1025+1402 and J1047+0739 both have H absorption, superimposed on the broad H emission. The H absorber of J1025+1402 has an EW of Å, and is blueshifted by km s-1 relative to the center of the broad H component. The H absorber in J1022+0841 can be described with two Gaussian components, one blueshifted by km s-1 and one redshifted by km s-1. Their EWs are Å and Å, respectively, yielding a total EW of Å. An alternative solution is one Gaussian absorber with km s-1 and EW of Å, but we find that this solution leaves more residuals as commented in Appendix B. J1047+0739 does not show clear Balmer absorption.
@@
For reference, we also present the Pa line profile in Figure 5. No absorption features associated with Pa are detected, consistent with the findings of Juodzbalis2024; Juodzbalis2025. This suggests that the Balmer absorbers originate from the neutral hydrogen elevated to the state but not significantly populated at .
All three targets show prominent blueshifted He I 10833 absorption. The He I 10833 absorber of J1025+1402 and J1047+0739 has multiple components with distinct kinematics (Figure 6). The line profile of the J1025+1402 He I 10833 absorber, as shown in Figure 5, can be described using three Gaussian components, blueshifted with , and km s-1, respectively. This profile may alternatively result from a mixture of saturated He I absorbers. The line profile of the J1047+0739 He I 10833 absorber can be described using five Gaussian components with , and km s-1, respectively. For J1022+0841, the He I 10833 absorber is well described by a single Gaussian component with km s-1. They highlight the prevalence of He I absorption in local LRDs. The He I line is a resonant transition from the metastable 23S state to the 23P state. The 23S state can be populated either through collisional excitation in high-density gas or via recombination from He+. It remains unclear whether the He I absorption in LRDs originates in the same dense clouds responsible for the Balmer absorption (Juodzbalis2024), where neutral He is collisionally excited into the 23S state, or instead in lower-density partially ionized outflows (Wang2025), where He is first ionized to He+ and then recombines into the 23S state. Although strong He I recombination lines (4473, 5877, 7067) with 23P as the lower state are detected in all three objects, the origin remains challenging to robustly constrain, as the emitting and absorbing gas clouds may not spatially co-exist. The complex kinematics of He I absorbers (Figure 6) may also be attributed to mixed origins.
| Name | H | H | He I 10833 | ||||||
| FWHM | EW | FWHM | EW | FWHM | EW | ||||
| J1025+1402 | 73 | ||||||||
| total | |||||||||
| J1047+0739 | No absorber | No absorber | |||||||
| total | |||||||||
| J1022+0841 | No absorber | ||||||||
| total | |||||||||